### 14 Mar 2026 v7

@@ 5.3.5 Linear attack analysis to $Eagle^{*}$ encryption algorithm
$Pr(C_{i}=c_{i}|P_{i}=p_{i})=\frac{1}{2^{L}}$
$Pr(Sx_{i}=sx_{i}|P_{i}=p_{i})=\frac{1}{2^{L}}$
$Pr(Sy_{i}=sy_{i}|P_{i}=p_{i})=\frac{1}{2^{L}}$
$Pr(S_{K+3}=s|P_{i}=p_{i})=\frac{1}{2^{L}}$
@@ 5.3.6 Differential attack analysis to $Eagle^{*}$ encryption algorithm
For the $Eagle^{*}$ encryption algorithm, suppose the differential attacker chooses two specific plaintexts $P_{1}$ and $P_{2}$, their difference is $\Delta$, that is $P_{2}=P_{1}+\Delta$, Since the ciphertexts in the second part and the third part are completely unrelated to the plaintext, only the ciphertext in the first part will be analyzed here, the corresponding ciphertexts are $C_{1}$ and $C_{2}$ , and the difference between the ciphertexts is $\varepsilon$ as $C_{2}=C_{1}+\varepsilon$. Since $C_{1}$ and $C_{2}$ are completely random, it is uncertain whether the difference $\varepsilon$ of the ciphertext is caused by randomness of $(x_{1},y_{1},x_{2},y_{2})$ or the uncertainty of the plaintext $(P_{1},P_{2})$. Furthermore, for any key $(\delta_{1},\delta_{2},\delta_{3})$, $x_{1}*\delta_{1}*\delta_{1}$ and $y_{1}*\delta_{2}*\delta_{2}$ subject to the uniform distribution, which can be denoted as
$Pr(C_{1}=c_{1},.,C_{K+3}=c_{K+3}|$
@@
$=1/2^{(K+3)L}$
That is to say, for any specific plaintext $P_{1}$ and $P_{2}$ selected by the attacker, after being encrypted with the same key, the corresponding block ciphertexts $(C_{1},C_{2},...,C_{K+3})$, $(Sx_{1},Sx_{2},...,Sx_{K+3})$ and $(Sy_{1},Sy_{2},...,Sy_{K+3})$ are completely random, and every possible value in the ciphertext space appears with equal probability. The attacker has no way to capture the propagation characteristics of the “difference” in the plaintext.

### 06 Apr 2026 v8

@@ 6.1.5 Linear attack analysis to $Eagle^{*}$ encryption algorithm
$Pr(C_{i}=c_{i}|P_{i}=p_{i})=\frac{1}{2^{L-1}}$
$Pr(Sx_{i}=sx_{i}|P_{i}=p_{i})=\frac{1}{2^{L-1}}$
$Pr(Sy_{i}=sy_{i}|P_{i}=p_{i})=\frac{1}{2^{L-1}}$
$Pr(Sz_{i}=sz_{i}|P_{i}=p_{i})=\frac{1}{2^{L}}$
$Pr(Su=su|P_{i}=p_{i})=\frac{1}{2^{L-2}}$
$Pr(P_{K+1}=p_{K+1}|P_{i}=p_{i})=\frac{1}{2^{L-1}}$
@@ 6.1.6 Differential attack analysis to $Eagle^{*}$ encryption algorithm
For the $Eagle^{*}$ encryption algorithm, suppose the differential attacker chooses two specific plaintexts $P_{1}$ and $P_{2}$, their difference is $\Delta$, that is $P_{2}=P_{1}+\Delta$, Since the ciphertexts in the second part and the third part are completely unrelated to the plaintext, only the ciphertext in the first part will be analyzed here, the corresponding ciphertexts are $C_{1}$ and $C_{2}$ , and the difference between the ciphertexts is $\varepsilon$ as $C_{2}=C_{1}+\varepsilon$. Since $C_{1}$ and $C_{2}$ are completely random, it is uncertain whether the difference $\varepsilon$ of the ciphertext is caused by randomness of $(x_{1},y_{1},x_{2},y_{2},u)$ or the uncertainty of the plaintext $(P_{1},P_{2})$. Furthermore, for any key $(\delta_{1},\delta_{2},\delta_{3})$, $x_{1}*\delta_{1}*\delta_{1}$, $y_{1}*\delta_{2}*\delta_{2}$ and $u*\delta_{3}*\delta_{3}$ subject to the uniform distribution, which can be denoted as
$Pr(C_{1}=c_{1},.,C_{K+1}=c_{K+1}|$
@@
$=1/2^{(K+1)(L-1)}$
$Pr(Sz_{1}=sz_{1},.,Sz_{K}=sz_{K}|$
$(P_{1}=p_{1},.,P_{K}=p_{K},W=(\delta_{1},\delta_{2},\delta_{3}))$
$=1/2^{KL}$
$Pr(Su=su,P_{K+1}=p_{K+1}|$
$(P_{1}=p_{1},.,P_{K}=p_{K},W=(\delta_{1},\delta_{2},\delta_{3}))$
$=1/2^{2(L-1)}$
That is to say, for any specific plaintext $P_{1}$ and $P_{2}$ selected by the attacker, after being encrypted with the same key, the corresponding block ciphertexts $(C_{1}$, $C_{2}$, …, $C_{K+1})$, $(Sx_{1}$, $Sx_{2}$, …, $Sx_{K+1})$, $(Sy_{1}$, $Sy_{2}$, …, $Sy_{K+1})$, $(Sz_{1}$, $Sz_{2}$, …, $Sz_{K})$, $Su,P_{K+1}$ are completely random, and every possible value in the ciphertext space appears with equal probability. The attacker has no way to capture the propagation characteristics of the “difference” in the plaintext.